OpenStudy (anonymous):

help medal fan Maritza is baking cookies to bring to school and share with her friends on her birthday.The recipe requires 3 eggs for every 2 cups of sugar.To have enough cookies for all of her friends,Maritza determined she would need 12 eggs.If her mom bought 6 cups of sugar,does martia have enough sugar to make cookies. why or why not.

OpenStudy (anonymous):

12/3=4 4*2=8 so no she doesn't have enough sugar. In this problem there is a 3:2 ratio. and since there are 12 eggs and 3*4=12 you need to multiply 2 by 4, which is 8. with only 6 cups of sugar she does not have enough sugar.

OpenStudy (anonymous):

You need two more cups of sugar dear

OpenStudy (anonymous):

i say @Ocicat_Kittens get a medal

OpenStudy (anonymous):

so what should i do for my work

OpenStudy (anonymous):

@,nmfrqew technically we are both right, but biggdaddymiller didn't fully answer the question, where I did, no offense biggdaddymiller

OpenStudy (anonymous):

could u tell me what should do for my work

OpenStudy (anonymous):

12/3=4 then 4*2=8 8>6 so she does not have enough sugar, that sound ok?

OpenStudy (anonymous):

it is awesome responce

OpenStudy (anonymous):

could u help me with a anoter queation i will posted

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

wait 5 min let me wtite this down please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm done with school for today so take your time

OpenStudy (anonymous):

k thanks

OpenStudy (leader):

@Ocicat_Kittens You answered it correctly :)

OpenStudy (leader):

Oh well it seems like @,nmfrqew tagged me but now deleted it.. lol

OpenStudy (anonymous):

i am going to pasted

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Oh no, I'm not here for the points, and you did spell it out. I'm just being lazy. Keep it up. I might need your brains!

OpenStudy (anonymous):

not to be mean @biggdaddymiller @Ocicat_Kittens help me very help

OpenStudy (anonymous):

Oh, no. I can't say never angry, but I always try to be polite, and pleasant.

OpenStudy (anonymous):

hope you don't mind, but I'm going to jump around answering questions, mention me when you're done